netfilter.org mirror setup HOWTO

Harald Welte laforge@netfilter.org $Revision: 4399 $

This document describes how to set up a FTP and/or WWW mirror of the netfilter.org project.

Contents

1 Introduction 2

2 Mirror of the www.netfilter.org homepage 2 2.1 Step zero: Do we need a mirror for your country? . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Step one: Prerequirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.2.1 rsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2.2 Webserver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2.3 Diskspace Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2.4 Traffic Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.3 Step two: Getting the netfilter homepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Step three: Configuring the name based virtual host . . . . . . . . . . . . . . . . . . . . . . . 3 2.5 Step four: Telling us about your new mirror . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Mirror of the ftp.netfilter.org 3

3.1 Which type of mirror? Setting up a mirror of the ftp server can be done in two ways. . . . . 3 3.1.1 Mirror as ftp.ccTLD.netfilter.org . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.2 Mirror as ftp.something.do.main/pub/netfilter . . . . . . . . . . . . . . . . . . . . . . 3

3.2 Step zero: Do we need a mirror for your country? . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.3 Step one: Prerequirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.3.1 rsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.3.2 Diskspace Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.3.3 Traffic Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.4 Step two: Getting the netfilter FTP tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.5 Step three: Telling us about your new mirror . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 The netfilter-mirrors mailinglist 4

5 Mirror of other netfilter.org services 5

1 Introduction

So you are interested in hosting a mirror of the netfilter.org project. This is good news, thank you for considering it. Please note that the netfilter/iptables project is quite popular, and the traffic to your website can be quite high. The netfilter.org main server kashyyyk.netfilter.org always has 100GB to 150GB per month, but that includes CVS and the mailinglists.

2 Mirror of the www.netfilter.org homepage

2.1 Step zero: Do we need a mirror for your country?

The netfilter/iptables project issues names for mirror sites depending on the country you are located. Usually a mirror would have a name like www.ccTLD.netfilter.org, where ccTLD is to be replaced with your country code top level domain (e.g. www.de.netfilter.org for germany).

Please make sure that we don’t already have a mirror for your country. At the moment we can only accept one mirror per country.

2.2 Step one: Prerequirements

2.2.1 rsync

The netfilter.org project offers the whole homepage directory via the rsync protocol. If you don’t know what rsync is, or don’t have rsync on your machine(s), we will not be able to use your mirror. Sorry.

More information about rsync can be found at the

rsync homepage http://samba.org/rsync/ .

2.2.2 Webserver

Apart from rsync, you will need a webserver that can server HTTP 1.1 name based virtual hosts.

2.2.3 Diskspace Requirements

The netfilter homepage is currently 50MB in size. I expect it to grow to 100, maybe 150MB in the future. Considering the size of today’s hard drives, I don’t think that anybody has problems with the size.

2.2.4 Traffic Requirements

Hard to tell. As we will not use any kind of load sharing, the distribution of the traffic happens ’manually’. This means, that the amount of traffic depends on the number of users that explicitly use www.ccTLD.netfilter.org instead of www.netfilter.org.

The total traffic of www.netfilter.org before we had any official mirrors was about 50GB per month.

2.3 Step two: Getting the netfilter homepage

You can download the netfilter homepage by using the following command:

     rsync -av --delete rsync.netfilter.org::www/ target_directory/

2.4 Step three: Configuring the name based virtual host

You will need to configure a name based virtual host for the following hostnames:

     www.ccTLD.netfilter.org
     www.ccTLD.iptables.org
     ccTLD.netfilter.org
     ccTLD.iptables.org
     mirror.netfilter.org
     mirror.iptables.org

2.5 Step four: Telling us about your new mirror

Send a short email with the IP address of your new homepage mirror to mirrormaster@netfilter.org mailto:mirrormaster@netfilter.org . We will then inform you once the DNS record has been entered into the netfilter.org and iptables.org zone file.

3 Mirror of the ftp.netfilter.org

3.1 Which type of mirror? Setting up a mirror of the ftp server can be done

in two ways.

3.1.1 Mirror as ftp.ccTLD.netfilter.org Since FTP doesn’t have a concept of name based virtual hosts, you would need to allocate an extra IP address for this ftp server. 3.1.2 Mirror as ftp.something.do.main/pub/netfilter This is the common case, where you mirror our FTP directory into a subdirectory of your FTP server. If you can put the mirror in the path given above (/pub/netfilter), we can also provide you with a ftp.ccTLD.netfilter.org hostname. If the directory appears somewhere else in your directory tree, we will not give you this kind of hostname for consistency reasons. Sorry.

3.2 Step zero: Do we need a mirror for your country?

The netfilter/iptables project issues names for mirror sites depending on the country you are located. Usually a mirror would have a name like ftp.ccTLD.netfilter.org, where ccTLD is to be replaced with your country code top level domain (e.g. ftp.de.netfilter.org for germany).

Please make sure that we don’t already have a mirror for your country. At the moment we can only accept one mirror per country.

3.3 Step one: Prerequirements

3.3.1 rsync

The netfilter.org project offers the whole homepage directory via the rsync protocol. If you don’t know what rsync is, or don’t have rsync on your machine(s), we will not be able to use your mirror. Sorry.

More information about rsync can be found at the

rsync homepage http://samba.org/rsync/ .

3.3.2 Diskspace Requirements

The netfilter FTP site is currently about 50MB in size. I expect it to grow to 150, maybe max. 250MB in the future. Considering the size of today’s hard drives, I don’t think that anybody has problems with the size.

3.3.3 Traffic Requirements

Hard to tell. As we will not use any kind of load sharing, the distribution of the traffic happens ’manually’. This means, that the amount of traffic depends on the number of users that explicitly use ftp.ccTLD.netfilter.org instead of ftp.netfilter.org.

The total traffic of ftp.netfilter.org before we had any official mirrors was about 50GB per month.

3.4 Step two: Getting the netfilter FTP tree

You can download the netfilter homepage by using the following command:

     rsync -av --delete rsync.netfilter.org::ftp/ target_directory/

3.5 Step three: Telling us about your new mirror

Send a short email with the IP address of your new FTP mirror to mirrormaster@netfilter.org mailto:mirrormaster@netfilter.org . We will then inform you once the DNS record has been entered into the netfilter.org and iptables.org zone file.

4 The netfilter-mirrors mailinglist

We have set up a mailinglist for discussion/announcements/questions/comments among the netfilter.org mirror administrators. If you want to host a mirror site, subscription to that list is mandatory. As soon as we put your IP address into our zone file, your email address will be subscribed to that mailinglist. Please note that posting to this list is restricted to members, so please don’t send mails with a ’From’ address different from the one subscribed. You can manage your list subscription at

the netfilter-mirrors@lists.netfilter.org admin interface http://lists.netfilter.org/mailman/listinfo/netfilter-mirrors

.

5 Mirror of other netfilter.org services

The netfilter/iptables project provides a number of other services, such as anonymous CVS, mailing list archives, bugzilla, CVSweb.

At this time we are not looking for mirrors for any of those services.